Sell or Process Further

St. Paul Lumber Company incurs a cost of $280 per hundred board feet in processing certain "rough-cut" lumber, which it sells for $320 per hundred board feet. An alternative is to produce a "finished cut" at a total processing cost of $350 per hundred board feet, which can be sold for $415 per hundred board feet.

a. What is the amount of the differential revenue?
$ per hundred board feet

b. What is the amount of the differential cost?
$ per hundred board feet

c. What is the amount of the differential income for processing rough-cut lumber into finished cut?
$ per hundred board feet
Answer #1

a) Differential revenue = 415-320 = 95 per hundred board feet

b) Differential cost = 350-280 = 70 per hundred board feet

c) Differential income = 95-70 = 25 per hundred board feet